### v1

# Excitation function of femtoscopic Lévy source parameters of pion pairs in EPOS4

Abstract
Three-dimensional (3D) femtoscopic source parameters of pions provide a sensitive probe of the space-time structure of particle-emitting sources in high-energy heavy-ion collisions. Compared to one-dimensional measurements, 3D femtoscopy not only provides a valuable cross-check but also offers a more complete characterization of the source geometry and its dynamical evolution. Particularly, differences between the "out" and "side" directions are sensitive to signals of a strong first-order phase transition, while the collision-energy dependence of Lévy radii may reveal non-monotonic features related to the equation of state (EoS).
In this work, we systematically investigate the transverse mass and collision-energy dependence of the three-dimensional femtoscopic parameters of pion pairs with Lévy-type sources in the STAR Beam Energy Scan (BES) range from $\sqrt{s_{NN}}=7.7$ to $200$ GeV using the EPOS4 model. The analyzed parameters include the Lévy index $\alpha$, the correlation strength $\lambda$, and the three-dimensional radii $R_{\mathrm{out}}$, $R_{\mathrm{side}}$ and $R_{\mathrm{long}}$ , derived quantities such as the radius difference $R_{\mathrm{diff}}=R_{\mathrm{out}}^{2}-R_{\mathrm{side}}^{2}$ and the ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$ are also investigated. The results show that the extracted radii $R_{\mathrm{side}}$ and $R_{\mathrm{long}}$ decrease with increasing transverse momentum and increase gradually with collision energy, while $R_{\mathrm{out}}$ shows little energy dependence. The Lévy index $\alpha$ exhibits only a mild dependence on $m_{T}$ and collision energy, whereas the correlation strength $\lambda$ shows a clear $m_{T}$ dependence and generally decreases with increasing collision energy. A comparison with EPOS3 results indicates general agreement within approximately $2\sigma$, with the notable exception of $R_{\mathrm{side}}$, which is systematically smaller in EPOS4.
*Corresponding author(s). E-mail(s):huang@ttk.elte.hu
1 Introduction
High-energy heavy-ion collisions provide a unique environment for studying strongly interacting matter under extreme conditions of temperature and density. One of the most sensitive tools for probing the space-time structure of the particle-emitting source is femtoscopy, which relies on the measurements of two-particle momentum correlations at small relative momenta.
Traditionally, the source distribution was modeled by a Gaussian form. However, in recent years, various experiments–from SPS [1, 2, 3], through RHIC [4, 5, 6, 7, 8], to LHC [9, 10]–have shown that Lévy-stable source distributions provide a better description of the measured correlation functions, particularly in cases where anomalous diffusion or long tails appear in the emission profile. Several physics mechanisms in high-energy heavy-ion collisions may contribute to such non-Gaussian behavior, including critical phenomena, jet fragmentation, and anomalous diffusion [11, 12, 13, 14]. The ELTE group has recently developed and applied a three-dimensional (3D) Lévy femtoscopy method to investigate pion sources in relativistic heavy-ion collisions [15]. This study provided a comprehensive framework for extracting the Lévy index of stability ($\alpha$), source radii ($R_{\rm{out}},R_{\rm{side}},R_{\rm{long}}$), and correlation strength ($\lambda$), offering improved sensitivity to non-Gaussian features of the source.
A particularly intriguing aspect of Lévy femtoscopy is its potential to reveal signatures of the QCD critical point. Near the critical point, long-range correlations are expected, leading to a non-trivial, possibly non-monotonic energy dependence of the Lévy index $\alpha$ [16]. Motivated by this, we extend our previous work by systematically analyzing the energy dependence of femtoscopic source parameters within the EPOS4 event generator framework. This provides a theoretical baseline for interpreting future experimental data and for exploring possible critical phenomena.
2 Methods
Our analysis is based on the latest 4.0.3 version of the general-purpose Monte Carlo event generator EPOS, to simulate heavy-ion collisions at different collision energies [17, 18]. Compared to EPOS3, EPOS4 implements a fully self-consistent, energy-momentum conserving parallel scattering scheme with subscattering dependent saturation scales. This represents a genuine paradigm shift that leads to harder high-multiplicity events and consequently modifies the collective source characteristics.
2.1 Distance distribution measurement and extraction of key parameters
In simulations, the pair distance distribution $D(\vec{\rho})$ can be directly obtained, where $\vec{\rho}$ represents the three-dimensional spatial separation vector of the particle pairs. This allows the extraction of key source parameters such as the Lévy index $\alpha$, the source size $R$, and the correlation strength $\lambda$. In experimental measurements, however, the spatial distributions are not directly accessible. Instead, one measures the correlation function in momentum space (as a function of momentum difference $\vec{q}$), which encodes information about both the spatial distribution $D(\vec{\rho})$ and the pair wave function $\psi_{\vec{q}}(\vec{\rho})$:
$C(\vec{q})=\int D(\vec{\rho})\,\bigl|\psi_{\vec{q}}(\vec{\rho})\bigr|^{2}\,d^{3}\vec{\rho}.$ (1)
For a Lévy-stable single-particle source, the pair distance distribution can be expressed as [19]:
$D(\vec{\rho})=\mathcal{L}(\alpha,2^{2/\alpha}{\cdot}R^{2};\vec{\rho})=\frac{1}{(2\pi)^{3}}\int d^{3}\vec{\zeta}\,e^{i\vec{\zeta}\cdot\vec{\rho}}\,e^{-\left|\vec{\zeta}^{\,T}R^{2}\vec{\zeta}\right|^{\alpha/2}},$ (2)
where $\vec{\zeta}$ is the integration variable, the superscript $T$ denotes the transpose, while $R^{2}=\mathrm{diag}\left(R_{\mathrm{out}}^{2},R_{\mathrm{side}}^{2},R_{\mathrm{long}}^{2}\right)$ denotes the diagonal matrix of Lévy-scale parameters defined in the Bertsch–Pratt “out-side-long” coordinate system [20, 21], neglecting cross-terms. The $2^{2/\alpha}$ factor appears due to $D$ being defined as the autoconvolution of the Lévy-distributed single-particle source.
In this study, for each energy, a total of 4000 minimum-bias events were generated using EPOS4 with its default X3F Equation of State with cross-over transition and three flavour conservation [22, 23]. Unlike Ref. [24], where the angle-averaged distance distribution $D(\rho)=\frac{1}{4\pi}\int D(\vec{\rho})\,d\Omega$ was measured, but similarly to Refs. [25, 15, 26], we investigate the 3D pair distribution. To express the source coordinates in the Bertsch-Pratt coordinates of the longitudinally co-moving system (LCMS) [27, 28, 29], the components of $\vec{\rho}$ are computed from the particle positions in the laboratory frame as follows:
$\rho_{\mathrm{out}}^{\mathrm{LCMS}}$ $=r_{x}\cos\varphi+r_{y}\sin\varphi-\frac{k_{T}}{K_{0}^{2}-K_{z}^{2}}\left(K_{0}t-K_{z}r_{z}\right),$ (3)
$\rho_{\mathrm{side}}^{\mathrm{LCMS}}$ $=-r_{x}\sin\varphi+r_{y}\cos\varphi,$ (4)
$\rho_{\mathrm{long}}^{\mathrm{LCMS}}$ $=\frac{K_{0}r_{z}-K_{z}t}{\sqrt{K_{0}^{2}-K_{z}^{2}}}.$ (5)
Here, the azimuthal angle $\varphi$ is defined by $\cos\varphi={K_{x}}/{k_{T}}$, $K_{0}$ and $(K_{x},K_{y},K_{z})$ denote the temporal and spatial components of the pair’s average four-momentum $\vec{K}$, $k_{T}=\sqrt{K_{x}^{2}+K_{y}^{2}}$ is the pair transverse momentum, and $(r_{x},r_{y},r_{z},t)$ are the particle coordinates in the laboratory frame. Detailed derivations and conventions can be found in the Methods sections of Refs. [15, 25].
For each event, the distribution of relative distances $D(\vec{\rho})$ between pion pairs was constructed in the longitudinally co-moving system. The distance distributions were then fitted using the Lévy parameterization, with the fitting procedure following the methodology outlined in Ref. [15], to extract the parameters $\alpha$, $R_{\rm{out}}$, $R_{\rm{side}}$, $R_{\rm{long}}$, and $\lambda$. The fits were found to describe the obtained sources well in all transverse mass bins at all collision energies. Thus, the interpretation of the extracted Lévy source parameters is adequate. As discussed in Ref. [25], the fitted distribution corresponds to the core-core source, consisting of pairs of pions produced close to the interaction point, i.e., primordial pions and the decay products of the shortest lived resonances.
2.2 Systematic uncertainties
The systematic uncertainties were estimated by varying several analysis conditions (sets) around the default configuration. The corresponding cuts applied for different sources are summarized below, with the numbers in parentheses indicating the values used for the various collision energies.
(1) Number of averaged events
Due to limited statistics in single-event analyses, events were merged to achieve sufficient pair statistics. It was carefully investigated and confirmed that the extracted parameters converge at a sufficiently large event number. For different collision energies, the number of merged events ($N_{\rm events}$) was varied as follows (the middle value is taken as default):
  • 7.7, 9.2, 11.5, and 14.5 GeV: (150, 170, 190)
  • 19.6, 27, 39, and 62.4 GeV: (120, 150, 180)
  • 130 and 200 GeV: (70, 100, 130)
(2) Pair momentum difference
In quantum-statistical correlations, the signal appears at limited values of relative momentum $Q_{\mathrm{LCMS}}$. In order to conceptually replicate experimental procedures, we utilize pairs up to a maximal momentum difference of $Q_{\mathrm{LCMS}}^{\mathrm{max}}=c\cdot\sqrt{m_{T}}({\rm MeV})$ where the coefficient $c$ was varied for different collision energies (again the middle value is the default):
  • 7.7, 9.2, 11.5, and 14.5 GeV: (10, 12, 15)
  • 19.6, 27, 39, 62.4, 130, and 200 GeV: (7, 10, 12)
(3) Fit range in $\rho$
The stability of the extracted source parameters was examined by varying the fitting range in $\rho$:
  • 7.7, 9.2, 11.5, and 14.5 GeV: (1-15, 1-20, 1-30)
  • 19.6, 27, 39, 62.4, 130, and 200 GeV: (1-20, 1-30, 1-40)
(4) Centrality definition
Centrality was defined using either the impact parameter (bimbim) from the EPOS4 definition or the charged-particle multiplicity ($N_{\mathrm{ch}}$). The two-dimensional ($bim,N_{\mathrm{ch}}$) distributions of the dataset were analyzed to define centrality bins according to the accumulated event fractions (0-10%, 10-20%, etc.). The centrality boundaries extracted from the impact parameter distribution are further found to be consistent with those provided directly by EPOS4. Overall, the choice of centrality definition contributes less than 2% to the systematic uncertainty of all extracted femtoscopic quantities.
[…]
$\sqrt{s_{{}_{NN}}}$ Source $\alpha$ $\lambda$ $R_{out}$ $R_{side}$ $R_{long}$
7.7 GeV $N_{\rm events}$ $0.59\%$ $0.25\%$ $0.24\%$ $0.28\%$ $0.45\%$
$Q_{\mathrm{LCMS}}^{\rm max}$ $0.66\%$ $0.13\%$ $0.17\%$ $0.98\%$ $2.15\%$
$\rho_{\rm fit}^{\rm min/max}$ $0.38\%$ $0.16\%$ $1.01\%$ $0.12\%$ $0.69\%$
Centrality $0.89\%$ $0.23\%$ $0.14\%$ $0.58\%$ $0.76\%$
Total $1.31\%$ $0.39\%$ $06\% 1.1$$18\%$ $2.43\%$
19.6 GeV $N_{\rm events}$ $0.30\%$ $14\% 0.0$$1\% 0.0$$09\%$ $0.16\%$
$Q_{\mathrm{LCMS}}^{\rm max}$ $48\% 0.0$$03\%$ $0.43\%$ $1.02\%$ $1.73\%$
$\rho_{\rm fit}^{\rm min/max}$ $1.34\%$ $0.36\%$ $43\% 0.0$$08\%$ $0.79\%$
Centrality $0.51\%$ $0.24\%$ $0.41\%$ $0.51\%$ $0.23\%$
Total $1.54\%$ $0.45\%$ $0.73\%$ $1.14\%$ $1.92\%$
200.0 GeV $N_{\rm events}$ $11\% 0.0$$06\%$ $13\% 0.0$$7\% 0.0$$06\%$
$Q_{\mathrm{LCMS}}^{\rm max}$ $2\% 0.440$$4\%$ $0.24\%$ $1.03\%$ $1.95\%$
$\rho_{\rm fit}^{\rm min/max}$ $0.53\%$ $0.14\%$ $0.93\%$ $0.27\%$ $0.85\%$
Centrality $\% 0.170.17\%$ $0.59\%$ $0.40\%$ $0.34\%$
Total $0.70\%$ $0.50\%$ $\% 1.141.14\%$ $2.16\%$
3 Results and discussions
In this section we present the results for the femtoscopic source parameters as functions of collision energy, focusing on three main aspects: source shape ($\alpha$), source size ($R_{o,s,l}$), and correlation strength ($\lambda$).
3.1 Source shape
The Lévy index $\alpha$ characterizes the shape of the emission source: $\alpha=2$ corresponds to a Gaussian source, while smaller values indicate long-tailed Lévy-stable distributions, often associated with contributions from long-lived resonances or anomalous diffusion-like dynamics.
[…]
Figure 1: Lévy index $\alpha$ from EPOS4 calculations as a function of transverse mass $m_{T}$ at different energies. The bands show the systematical uncertainties.
Figure 2: Lévy index $\alpha$ from EPOS4 calculations as a function of collision energy $\sqrt{s_{NN}}$ in three bins of $k_{T}$. The bands show the systematical uncertainties.
3.2 Source size
The extracted source radii exhibit the expected decrease with increasing transverse momentum and an overall growth with increasing collision energy, reflecting the expansion dynamics of the system. The energy dependence of $R_{o}$, $R_{s}$, and $R_{l}$ provides a systematic baseline for comparison with future experimental femtoscopy results.
[…]
Figure 3: Lévy radii from EPOS4 calculations as a function of transverse momentum $m_{T}$ at $\sqrt{s_{NN}}=7.7-200\penalty 10000\ GeV$. The bands show the systematical uncertainties.
Figure 4: Energy dependence of Lévy radii from EPOS4 calculations at $k_{T}=0.375-0.425$ GeV/$c$. The bands show the systematical uncertainties.
Ref. [16] has highlighted the sensitivity of the Lévy scale difference $R_{\mathrm{diff}}^{2}=R_{\mathrm{out}}^{2}-R_{\mathrm{side}}^{2}$ to potential critical-point signatures in the QCD phase diagram. Motivated by this, we investigated the energy dependence of both $R_{\mathrm{diff}}$ and the ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$, as shown in Figs. 6 and 6. The quantity $R_{\mathrm{diff}}$ decreases with increasing $m_{T}$, consistent with the behavior of the individual femtoscopic radii. And its collision-energy dependence shows a non-monotonic pattern, exhibiting a noticeable drop at 14.5 GeV, but without a clear global trend across the full energy range. The ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$, on the other hand, exhibits a stonger overall decrease with increasing collision energy, accompanied by a notable pronounced enhancement in the interval 14.53914.5-39 GeV. The investigation of these observables for different equations of state is of crucial importance and presents the focus of subsequent studies.
Figure 5: The radius difference $R_{\mathrm{diff}}^{2}$ from EPOS4 calculations as a function of collision energy $\sqrt{s_{NN}}$ in three $k_{T}$ bins. The bands show the systematical uncertainties.
Figure 6: The radius ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$ from EPOS4 calculations as a function of collision energy $\sqrt{s_{NN}}$ at $k_{T}$=0.375-0.525GeV/cGeV/c. The bands show the systematical uncertainties.
3.3 Correlation strength
The correlation strength parameter $\lambda$ is sensitive [31, 32] to the fraction of coherent emission, long-lived resonance contributions, and purity corrections. In simulations, $\lambda$ can be extracted from the integral of the fitted core-core component of the source [25]. Values extracted in this EPOS4 analysis show a moderate decrease with increasing collision energy, which may be attributed to the enhanced role of resonance decays at higher energies.
The transverse-mass dependence of the extracted $\lambda$ parameter at various collision energies is presented in Fig. 8, showing an overall increase with $m_{T}$. As shown in Fig. 8, $\lambda$ reveals a decreasing trend with increasing $\sqrt{s_{NN}}$, although noticeable anomalies are observed at $\sqrt{s_{NN}}=9.2$ and $11.5$ GeV, where the values deviate from the general behavior.
Figure 7: Correlation strength parameters $\lambda$ from EPOS4 as a function of transverse momentum $m_{T}$ at $\sqrt{s_{NN}}=7.7-200\penalty 10000\ GeV$. The bands show the systematical uncertainties.
Figure 8: Energy dependence of $\lambda$ from EPOS4 at $k_{T}=0.375-0.525\penalty 10000\ GeV/c$. The bands show the systematical uncertainties.
3.4 Comparison to data and EPOS3
In this analysis, the femtoscopic parameters of pions from $200$ GeV Au+Au collisions are generally consistent with those obtained from EPOS3 in Ref. [25] within approximately $2\sigma$, with the notable exception of the $R_{\mathrm{side}}$ radius, which exhibits a systematic reduction. For $R_{\mathrm{side}}$, EPOS4 predicts systematically smaller values, with deviations corresponding to $N_{\sigma}=2.5$$5.8$, where $N_{\sigma}={\Delta}/{\sigma_{\mathrm{combined}}}$ and $\Delta$ is the difference between the EPOS4 and EPOS3 results, while $\sigma_{\mathrm{combined}}=\sqrt{\sigma_{\mathrm{EPOS4}}^{2}+\sigma_{\mathrm{EPOS3}}^{2}}$ denotes the quadrature sum of their uncertainties. Due to this similarity to EPOS3, and the detailed data comparisions of EPOS3 to data of Ref. [7] in Ref. [25], it can be stated that EPOS4 is also compatible with the data.
Due to its particular importance [7], we nevertheless discuss the data comparision in case of the $\lambda$ parameter. As mentioned, Ref. [25] provided a detailed discussion showing that the experimental intercept parameter $\lambda$ can be regarded as the integral of $D_{\mathrm{cc}}$, which represents the core-core component of the particle-emitting source, when the total number of pairs is normalized to unity. It is important to note however, that pairs at large distances consist of pions produced far away from the interaction point, thus their “distance of closest approach” would be sizeable—resulting in its removal from the actual pion sample in an experimental analysis. Thus in fact it is the number of pion pairs up to a given distance that constitute the “full” pion sample. In order to make a direct comparison with the experimental results, similarly to Ref. [25], different integration limits $\rho_{\lambda}^{max}$ (0.1 mm, 1 cm, and 50 cm) were applied in this work to obtain $\lambda$, which was then compared with the PHENIX experimental data at $\sqrt{s_{NN}}=200$ GeV [7]. The $m_{\mathrm{T}}$ dependences of $\lambda$ and the scaled $\lambda/\lambda_{\rm max}$ are shown in Fig. 10 and Fig. 10, respectively. While the nominal $\lambda$ value can be described only if choosing a small $\rho_{\lambda}^{max}$, the description of the scaled correlation strength measurements by EPOS4 is adequate in all cases.
Figure 9: Correlation strength parameters $\lambda$ from EPOS4 for variotus choices of $\rho_{\lambda}^{max}$ (lines of different colors with bands), compared to PHENIX data of Ref. [7] (black points), as a function of mT. The bands shows the systematical uncertainties of the EPOS4 calculations. The statistical and systematical uncertainties of the PHENIX data points are shown as error bars and boxes, respectively.
Figure 10: Normalized correlation strength parameters $\lambda$$\lambda_{max}$ from EPOS4 for various choices of $\rho_{\lambda}^{max}$ (lines of different colors with bands), compared to PHENIX data of Ref. [7] (black points), as a function of mT. The bands show the systematical uncertainties of the EPOS4 calculations. The statistical and systematical uncertainties of the PHENIX data points are shown as error bars and boxes, respectively.
4 Summary and outlook
We presented detailed investigations of the simulated particle emitting source in $\sqrt{s_{NN}}=7.7-200$ GeV 0-10% centrality Au+Au collisions within the framework of the EPOS4 model. We calculated the spatial pair distribution $D(\vec{\rho})$ for identical charged pions in the longitudinally comoving system, for several bins of pair transverse mass ($m_{T}$) and calculated their one-dimensional projections in the Bertsch-Pratt out, side, and long coordinates. We fitted these projections simultaneously with projections of a single 3D Lévy distribution, and found a statistically and qualitatively also adequate description of the source, up to distances of a few dozen (or up to a hundred) femtometers. We then extracted the following source parameters: Lévy index (or shape parameter) $\alpha$, spatial Lévy scales (or HBT radii) $R_{\rm out,side,long}$, and correlation strength parameter $\lambda$. We then investigated the $m_{T}$ and $\sqrt{s_{NN}}$ dependence of these source parameters.
We found that the radii $R_{\rm out,side,long}$ decrease with transverse mass at all collision energies, a characteristic observation attributed to the expansion of the system. This decrease appeared to be the strongest in the longitudinal direction. The radii show a moderate increase with collision energy, also strongest for the longitudinal direction. An average scale ($R_{\rm avg}$) and out-side anisotropy (in form of the difference and the ratio of the out and side radii) were also investigated. The anisotropy appears to decrease towards higher collision energies. One of the important future research directions is to investigate this for a modified equation of state, as the change in the order of the phase transition is expected to have characteristic signature in the increase of the correlation radius in the out direction ($R_{\rm out}$).
The Lévy exponent, describing the source shape, was also investigated as a function of transverse mass and collision energy. A weak but smooth increase with transverse mass is apparent for all collision energies. A less smooth decrease is found as a function of collision energy. As the $\alpha$ parameter is expected to be strongly affected by the equation of state, in particular a decrease in $\alpha$ is expected at the critical point, again it is one of the crucial research directions to calculate $\alpha$ versus $\sqrt{s_{NN}}$ for an equation of state describing a second order phase transition.
The correlation strength parameter $\lambda$, extracted based on the integral of the fitted Lévy distribution, and corresponding to the core fraction among all pion pairs, was also investigated as a function of transverse mass and collision energy. A clear increase is found with transverse mass at low $m_{T}$, for all collision energies. This can also be described as a “hole” at small $m_{T}$, as discussed by PHENIX in Ref. [7]. EPOS4, similarly to EPOS3 [25], describes the scaled correlation strength $\lambda/\lambda_{\rm max}$ PHENIX data well. Furthermore, a weak decrease is found in $\lambda$ for larger collision energies.
It is important to investigate the equation of state dependence of these observables, and compare them to data, once Lévy source parameter measurements at several collision energies will be available.
5 Acknowlegements
This research was funded by the NKFIH grants TKP2021-NKTA-64, PD-146589, K-146913, K-138136, and NKKP-152097.